# LLaVA-SP: Enhancing Visual Representation with Visual Spatial Tokens for MLLMs

### v1

@@ 1 Introduction
To address this question, we propose LLaVA-SP to enhance the visual representation of MLLMs. The Projector of LLaVA-SP consists of two key designs: the Spatial Feature Extractor (SFE) and the Detail Feature Integrator (DFI). 1) The SFE aims to enhances the feature representation of the vision encoder by adding only six visual spatial tokens. These six visual spatial tokens can be introduced through two operations: cropping or pooling. The motivation for cropping is to emphasize detailed regional features, while pooling captures the image’s overall information. In the cropping approach, we progressively crop the ViT patch features inward until reaching the central region, obtaining multi-scale features. These features are then arranged from left to right in the order of “from central region to global”. Cropping focuses on regional detail, making it suitable for tasks requiring fine-grained image understanding. In contrast, the pooling method uses adaptive pooling layers to generate multi-scale features that capture varying levels of abstraction, which are then arranged from left to right in the order of “from abstract to specific”. This strategy is inspired by the hierarchical manner in which humans perceive or create images [52], first capturing the global structure and then focusing on local details. Pooling is especially beneficial for tasks that require a more general understanding of the image. For both methods, the ViT patch features are reshaped to their original 2D shapes. They are then reorganized according to either the “from central region to global” or “from abstract to specific” strategy, resulting in structured multi-scale features. Finally, convolutional kernels of varying sizes are applied to these multi-scale features to capture visual spatial tokens, which are then concatenated with the original visual tokens to form a comprehensive visual representation. 2) The DFI further enhances visual spatial features through a cross-attention mechanism. Without increasing the number of visual spatial tokens extracted by SFE, DFI derives fine-grained features from the large-size visual feature maps and integrates them into the visual spatial tokens to accomplish feature fusion, which further enhances the visual representation and thereby improves the detailed understanding ability of MLLMs.
@@ 2.1 Multimodal Large Language Models
Research in multimodal large language models has focused on aligning visual and linguistic representation to improve interaction between the two domains. Flamingo [1] introduced the Perceiver Resampler, which employed a cross-attention mechanism to integrate visual data into large language models (LLMs). The BLIP [26, 27, 57, 12] and Qwen-VL [2, 54] series developed the Q-former structure for visual-language alignment, using learnable parameter queries to compress visual information and reduce the number of visual tokens. Alternatively, the Mini-GPT4 [67] and LLaVA series [34, 33] adopted a simple multilayer perceptron (MLP) as projectors to map visual features into the language representation space of LLMs.
@@
End-to-end MLLMs represent a cutting-edge area of research, focusing on direct processing of visual inputs without relying on pre-trained vision encoder. Fuyu-8B [45], EVE [13], SOLO [7], and OtterHD [25], forgo pre-trained vision encoder and directly segment images into patches for input into LLMs instead. These methods allow MLLMs to bypass the limitations imposed by the prior knowledge of vision encoder, facilitating the learning of unaltered visual information. Our work builds on LLaVA-1.5, investigating the potential of vision encoder to enhance visual representation for MLLMs.
@@ 4.6 Methods Generalization
We replaced CLIP-ViT-L/14-336 with SigLIP-L/16-384 and applied SP method to InternVL-2.0. Our method focuses on enhancing the visual representation of CLIP, effectively adding an external module to CLIP. Other MLLMs, which involve higher resolutions, more visual tokens, and stronger vison encoder, are orthogonal to our approach, as their CLIP still has representational limitations. Tab. 7 demonstrates that our approach can be adapted to stronger vision encoder and the novel MLLM framework.

### v2

@@ 1 Introduction
To address this question, we propose LLaVA-SP to enhance the visual representation of MLLMs. The Projector of LLaVA-SP consists of two key designs: the Spatial Feature Extractor (SFE) and the Detail Feature Integrator (DFI). 1) The SFE aims to enhance the feature representation of the vision encoder by adding only six visual spatial tokens. These six visual spatial tokens can be introduced through two operations: cropping or pooling. The motivation for cropping is to emphasize detailed regional features, while pooling captures the image’s overall information. In the cropping approach, we progressively crop the ViT patch features inward until reaching the central region, obtaining multi-scale features. These features are then arranged from left to right in the order of “from central region to global”. Cropping focuses on regional details, making it suitable for tasks requiring fine-grained image understanding. In contrast, the pooling method uses adaptive pooling layers to generate multi-scale features that capture varying levels of abstraction, which are then arranged from left to right in the order of “from abstract to specific”. This strategy is inspired by the hierarchical manner in which humans perceive or create images [52], first capturing the global structure and then focusing on local details. Pooling is especially beneficial for tasks that require a more general understanding of the image. For both methods, the ViT patch features are reshaped to their original 2D shapes. They are then reorganized according to the “from central region to global” or “from abstract to specific” strategy, resulting in structured multi-scale features. Finally, convolutional kernels of varying sizes are applied to these multi-scale features to capture visual spatial tokens, which are then concatenated with the original visual tokens to form a comprehensive visual representation. 2) The DFI further enhances visual spatial features through a cross-attention mechanism. Without increasing the number of visual spatial tokens extracted by SFE, DFI derives fine-grained features from the large-size visual feature maps and integrates them into the visual spatial tokens to accomplish feature fusion, which further enhances the visual representation and thereby improves the detailed understanding ability of MLLMs.
@@ 2.1 Multimodal Large Language Models
Research in multimodal large language models has focused on aligning visual and linguistic representation to improve interaction between the two domains. Flamingo [1] introduced the Perceiver Resampler, which employed a cross-attention mechanism to integrate visual data into large language models (LLMs). The BLIP [26, 27, 57, 12] and Qwen-VL [2, 54] series developed the Q-former structure for visual-language alignment, using learnable parameter queries to compress visual information and reduce the number of visual tokens. Alternatively, the Mini-GPT4 [67] and LLaVA series [34, 33] adopted a simple multilayer perceptron (MLP) as projector to map visual features into the language representation space of LLMs.
@@
End-to-end MLLMs represent a cutting-edge area of research, focusing on direct processing of visual inputs without relying on pre-trained vision encoder. Fuyu-8B [45], EVE [13], SOLO [7], and OtterHD [25] forgo pre-trained vision encoder and directly segment images into patches for input into LLMs instead. These methods allow MLLMs to bypass the limitations imposed by the prior knowledge of vision encoder, facilitating the learning of unaltered visual information. Our work builds on LLaVA-1.5, investigating the potential of vision encoder to enhance visual representation for MLLMs.
@@ 4.6 Methods Generalization
We replaced CLIP-ViT-L/14-336 with SigLIP-L/16-384 and applied SP method to InternVL-2.0. Our method focuses on enhancing the visual representation of CLIP, effectively adding an external module to CLIP. Other MLLMs, which involve higher resolutions, more visual tokens, and stronger vision encoder, are orthogonal to our approach, as their CLIP still has representational limitations. Tab. 7 demonstrates that our approach can be adapted to stronger vision encoder and the novel MLLM framework.